# XXLTraffic: Expanding and Extremely Long Traffic forecasting beyond test adaptation

### v2

@@ 5. Experiments
We conducted experiments for both extremely long forecasting with gaps using gap dataset and conventional long-term forecasting using hourly dataset and daily dataset. Additionally, referring to the definition in Fig. 3, we set the gap parameter $g$ as 1 year, 1.5 years, and 2 years for the gap dataset, as illustrated by Fig. 6.
@@ 5.1. Datasets
We conducted experiments on all proposed sub-datasets. To maintain consistency with previous state-of-the-art benchmarks, we selected districts 03, 04, and 08 (widely recognized as PEMS03/04/08) for the experiments using the gap, and districts 03, 04,07 and 08 for hourly, and daily experiments. Results for other datasets are presented in Appendix. All sub-datasets were divided into training, validation, and test sets using a 6:2:2 ratio. For the gap dataset, due to the extensive span of up to 20 years resulting in a large sample size, we fixed a seed during data preprocessing to select 10% of the dataset for training and testing to quickly demonstrate our results. The details of the datasets used in our benchmarking is in Appendix A.1.
Table 3. Comparison in gap dataset. The bold text indicates the best.
@@ 5.2. Baselines
In our comparison experiments, we adopted four popular baselines, including MLP, Transformer, and Mamba architectures. Informer (Zhou et al., 2021) introduces an efficient transformer for long sequence time-series forecasting using ProbSparse self-attention and self-attention distilling, enabling encoder-decoder architectures to handle long sequences effectively. MICN (Wang et al., 2023a) proposes a multi-scale context network that models both local and global contexts for long-term time series forecasting, capturing patterns across different temporal scales to enhance performance. FEDformer (Zhou et al., 2022) introduces a frequency-enhanced decomposed transformer that models time series in both time and frequency domains, improving long-term forecasting by effectively capturing temporal patterns. PatchTST (Nie et al., [n. d.]) applies transformers to time series by treating them as sequences of patches, enabling effective long-term forecasting through self-attention over patch representations to capture temporal dependencies.Autoformer (Wu et al., 2021), an earlier state-of-the-art model, leverages a decomposition architecture and auto-correlation mechanism to enhance efficiency and accuracy in long-term time series forecasting, outperforming traditional Transformer models. iTransformer (Liu et al., 2024a) is the latest and most effective Transformer-based model, utilizing attention and feed-forward networks on inverted dimensions, embedding time points into variate tokens. DLinear (Zeng et al., 2023) challenges the effectiveness of Transformer models by proposing a simple one-layer linear model that captures temporal relations in an ordered set of continuous points. It employs positional encoding and uses tokens to embed sub-series, preserving some ordering information in Transformers. Lastly, Mamba (Gu and Dao, 2023), a well-known sequential model from last year, uses a bidirectional Mamba block to extract inter-variate correlations and temporal dependencies. Additionally, we have selected five SOTA baselines  (Yu et al., 2018; Guo et al., 2019; Wu et al., 2019; Bai et al., 2020; Jiang et al., 2023) from traffic forecasting domain.
@@ 5.3. Implementation Details
We adopted the default settings of the Time-Series-Library (Wu et al., 2022) to conduct a comprehensive comparison of baselines. We use the results of five random seeds as the average. We used 96 time steps as input and 336 time steps as ground truth. The code was implemented in PyTorch and executed on a V100 GPU with 32GB memory and 384GB RAM, provided by NCI Australia, an NCRIS-enabled capability supported by the Australian Government.
@@ 5.4. Results of Extremely Long Forecasting with gaps
We use Mean Squared Error (MSE) and Mean Absolute Error (MAE) metrics to evaluate performance, averaging results across different seeds. It is observed that nearly all results are poor, highlighting the significant challenge posed by domain shifts over time for extremely long forecasting with gaps. These baseline results also indicate that traditional the-state-of-the-art (SOTA) rankings and methodologies are no longer effective. Notably, MICN, which performs the worst under conventional data and settings, shows the best performance in our setting. This is understandable because MICN’s core technology focuses on exploring the correlation within the data itself. This insight suggests that future efforts to tackle this problem should place greater emphasis on leveraging the intrinsic potential of the data. Additionally, we have compiled the training time for one epoch of all the baselines on the largest dataset, PEMS04_gap, and the smallest dataset, tfNSW, as shown in Table 9.
Considering that our dataset provides an extensive temporal range for training, we can theoretically extend the input step length considerably. We extended the 96 steps input from Table 3 to a maximum of 1440 steps for testing. As shown in Table 4, the performance improves significantly with the increase in input step length, further demonstrating the substantial potential of our dataset for deep exploration.
@@ 5.5. Results of Hourly and Daily Forecasting
We believe that both the hourly and daily datasets are equally significant. Multi-scale, diverse datasets can provide the community with valuable references. We observe that the performance degrades progressively from the hourly to the daily to the gap datasets. Smaller time scales help reduce complexity and uncertainty, thereby improving the accuracy of prediction. Research has shown that clustering at different scales can improve model performance (Wang et al., 2024). Therefore, our aggregated version of the data will contribute new external features to the community.
@@
6. Prospects and Constraints
Prospects. Our dataset spans the longest time period among existing datasets, and it is not only the largest spatially, but also evolving in growth. It can continue to update according to the updates from the PeMS system in the future. It is specifically designed for various complex scenarios, such as those already mentioned, including extremely long forecasting with long gaps, and hourly and daily predictions.
Constraints. The limitations of our dataset are also quite evident. Due to the sheer size of our dataset, it requires more computational resources. However, with the maturation of large language models and foundational models, we believe its large volume will become an advantage, contributing more diverse data to the spatio-temporal large model community.
Appendix A Appendix
A.1. Dataset
To support our setting of extremely long forecasting with gaps, we selected a subset installed in the early stages and have consistently collected new data up to the present, which is shown as follows:
Datasets(Gap/Hour/Day) Time Period Nodes
PEMS03_gap&agg 03/2001 - 03/2024 151
PEMS04_gap&agg 06/2002 - 03/2024 822
PEMS05_gap&agg 03/2012 - 03/2024 103
PEMS06_gap&agg 12/2009 - 03/2024 130
PEMS07_gap&agg 06/2002 - 03/2024 3062
PEMS08_gap&agg 03/2001 - 03/2024 212
PEMS10_gap&agg 06/2007 - 03/2024 107
PEMS11_gap&agg 09/2002 - 03/2024 521
PEMS12_gap&agg 01/2002 - 03/2024 1543
tfNSW 01/2013 - 05/2024 27
A.2. Results
Here we provided the results of District 5,6,10,11,12, as shown in Table 7 and Table 8:
Table 7. Comparison in hourly and daily datasets in District 05,06,10,11,12. The symbol -- indicates that the result is an outlier.
@@ A.2. Results
Table 8. Comparison in gap dataset in District 05,06,07,10,11,12. The bold text indicates the best.
Methods Mamba iTransformer DLinear Autoformer
Metrics MSE MAE MSE MAE MSE MAE MSE MAE
PEMS05_gap 1-year gap 96 2.079 1.209 1.945 1.164 1.291 0.916 1.065 0.796
192 2.132 1.256 1.984 1.185 1.750 1.099 1.063 0.809
336 2.377 1.340 2.067 1.234 1.894 1.144 1.135 0.827
1.5-year gap 96 1.852 1.122 1.879 1.078 1.683 1.054 1.060 0.785
192 1.929 1.182 1.593 1.032 1.633 1.045 0.912 0.712
336 2.370 1.313 2.214 1.071 0.794 1.184 1.345 0.882
2-year gap 96 1.868 1.106 1.580 0.969 1.602 1.018 0.828 0.672
192 2.219 1.274 1.481 0.958 1.589 1.027 1.018 0.772
336 2.695 1.212 2.207 1.201 1.922 1.139 1.186 0.839
PEMS06_gap 1-year gap 96 1.806 1.066 0.875 0.692 1.173 0.837 1.216 0.859
192 1.928 1.112 1.227 0.848 1.410 0.942 0.961 0.751
336 2.181 1.212 1.594 1.003 1.501 0.976 0.992 0.769
1.5-year gap 96 1.549 0.997 1.331 0.891 1.484 0.963 0.885 0.710
192 1.746 1.054 1.077 0.778 1.353 0.920 1.010 0.768
336 1.605 1.018 1.500 0.961 1.587 1.011 0.955 0.739
2-year gap 96 1.226 0.851 1.864 1.106 1.691 1.033 1.013 0.768
192 0.949 0.720 1.343 0.879 1.259 0.858 0.853 0.691
336 0.945 0.710 1.550 0.970 1.415 0.934 0.955 0.739
PEMS07_gap 1-year gap 96 1.719 1.006 1.756 1.024 1.680 1.024 1.215 0.867
192 1.637 0.996 1.816 1.061 1.776 1.043 1.202 0.852
336 1.637 0.991 1.784 1.028 1.774 1.034 1.098 0.774
1.5-year gap 96 1.571 0.972 1.705 1.016 1.585 0.987 1.184 0.844
192 1.578 0.982 1.657 1.004 1.624 0.996 1.088 0.800
336 1.818 1.127 1.712 1.028 1.668 1.036 0.991 0.729
2-year gap 96 5.411 1.319 2.055 1.149 1.746 1.048 1.099 0.802
192 12.620 1.370 2.006 1.128 1.705 1.002 1.284 0.875
336 9.614 1.378 1.812 1.057 1.605 0.968 0.972 0.734
PEMS10_gap 1-year gap 96 2.310 1.233 0.882 0.698 1.466 0.963 1.208 0.865
192 2.878 1.396 1.606 0.993 1.834 1.106 1.152 0.860
336 2.584 1.327 1.413 0.825 1.887 1.121 1.130 0.830
1.5-year gap 96 2.181 1.206 1.812 1.052 1.765 1.057 1.368 0.913
192 2.525 1.318 1.645 0.986 1.791 1.069 1.151 0.828
336 2.488 1.310 2.051 1.134 1.976 1.139 1.111 0.809
2-year gap 96 1.165 0.822 2.772 1.383 1.977 1.119 1.188 0.845
192 1.082 0.792 1.974 1.084 1.490 0.932 0.971 0.751
336 1.021 0.764 2.157 1.168 1.619 0.989 1.096 0.809
PEMS11_gap 1-year gap 96 5.199 0.891 5.417 0.997 5.276 0.936 4.930 0.854
192 5.300 0.920 5.504 1.029 5.393 0.974 5.114 0.867
336 5.251 0.901 5.410 0.985 5.364 0.955 4.830 0.849
1.5-year gap 96 5.871 0.968 6.045 1.296 5.914 1.224 5.691 0.878
192 5.968 1.012 6.121 1.314 5.993 1.229 5.792 0.884
336 6.167 1.074 6.214 1.326 6.025 1.299 5.947 0.967
2-year gap 96 5.914 0.996 6.136 1.318 5.945 1.243 5.761 0.978
192 6.541 1.043 6.213 1.327 6.014 1.289 6.245 1.001
336 6.541 1.086 6.221 1.332 6.024 1.300 6.268 1.024
PEMS12_gap 1-year gap 96 1.751 1.025 1.624 1.002 1.611 1.005 1.060 0.789
192 1.726 1.024 1.424 0.929 1.537 0.972 1.150 0.840
336 1.751 1.029 1.672 1.015 1.683 1.017 0.889 0.719
1.5-year gap 96 1.554 0.967 1.479 0.921 1.468 0.910 0.954 0.875
192 1.401 0.898 1.314 0.867 1.301 0.862 0.943 0.869
336 1.417 0.906 1.322 0.869 1.298 0.859 0.921 0.846
2-year gap 96 0.956 0.704 0.876 0.671 0.872 0.664 0.846 0.659
192 0.846 0.656 0.813 0.653 0.806 0.649 0.785 0.628
336 0.814 0.628 0.789 0.631 0.776 0.624 0.754 0.617
[…]
Table 9. Model Training Time Comparison. The training time for all baselines per epoch is measured in seconds.
Baselines Mamba iTransformer DLinear Autoformer Informer FEDFormer MICN PatchTST STGCN ASTGCN GWN AGCRN PDFormer
PEMS04_gap1 396.3 1310.2 237.9 1656.7 652.5 900.9 409.8 3649.4 2425.9 7921.9 4110.4 13369.1 8013.9
tfNSW_gap1 13.7 44.1 7.4 148.2 73.3 380.1 101.2 55.3 37.3 228.8 50.4 658.6 47.9

### v3

@@ 5. Experiments
We conducted experiments for both extremely long forecasting with gaps using gap dataset and conventional long-term forecasting using hourly dataset and daily dataset. Additionally, referring to the definition in Fig 5, we set the gap parameter $g$ as 1 year, 1.5 years, and 2 years for the gap dataset, as illustrated by Fig 5.
@@ 5.1. Datasets
We conducted experiments on all proposed sub-datasets. To maintain consistency with previous state-of-the-art (SOTA) benchmarks, we selected districts 03, 04, and 08 (widely recognized as PEMS03/04/08) for the experiments using the gap which is shown in Table 3. Results for other gaps and aggregated datasets on some baselines are presented in Tables˜4, 8 and 9. All sub-datasets were divided into training, validation, and test sets using a 6:2:2 ratio. For the gap dataset, due to the extensive span of up to 20 years resulting in a large sample size, we fixed a seed during data preprocessing to select 10% of the dataset for training and testing to quickly demonstrate our results. Details of the datasets used in our benchmarking is in Table 5.
Table 3. Comparison in gap dataset. Bold numbers indicate the best results.
@@
Table 4. Comparison in datasets in District 05,06,07,10,11,12. Bold numbers indicate the best results.
Datasets PEMS05 PEMS06
Metrics 1-year 1.5-year 2-year 1-year 1.5-year 2-year
Mamba 96 MSE/MAE 2.079 / 1.209 1.852 / 1.122 1.868 / 1.106 1.806 / 1.066 1.549 / 0.997 1.226 / 0.851
192 MSE/MAE 2.132 / 1.256 1.929 / 1.182 2.219 / 1.274 1.928 / 1.112 1.746 / 1.054 0.949 / 0.720
336 MSE/MAE 2.377 / 1.340 2.370 / 1.313 2.695 / 1.212 2.181 / 1.212 1.605 / 1.018 0.945 / 0.710
iTransformer 96 MSE/MAE 1.945 / 1.164 1.879 / 1.078 1.580 / 0.969 0.875 / 0.692 1.331 / 0.891 1.864 / 1.106
192 MSE/MAE 1.984 / 1.185 1.593 / 1.032 1.481 / 0.958 1.227 / 0.848 1.077 / 0.778 1.343 / 0.879
336 MSE/MAE 2.067 / 1.234 2.214 / 1.071 2.207 / 1.201 1.594 / 1.003 1.500 / 0.961 1.550 / 0.970
DLinear 96 MSE/MAE 1.291 / 0.916 1.683 / 1.054 1.602 / 1.018 1.173 / 0.837 1.484 / 0.963 1.691 / 1.033
192 MSE/MAE 1.750 / 1.099 1.633 / 1.045 1.589 / 1.027 1.410 / 0.942 1.353 / 0.920 1.259 / 0.858
336 MSE/MAE 1.894 / 1.144 0.794 / 1.184 1.922 / 1.139 1.501 / 0.976 1.587 / 1.011 1.415 / 0.934
Autoformer 96 MSE/MAE 1.065 / 0.796 1.060 / 0.785 0.828 / 0.672 1.216 / 0.859 0.885 / 0.710 1.013 / 0.768
192 MSE/MAE 1.063 / 0.809 0.912 / 0.712 1.018 / 0.772 0.961 / 0.751 1.010 / 0.768 0.853 / 0.691
336 MSE/MAE 1.135 / 0.827 1.345 / 0.882 1.186 / 0.839 0.992 / 0.769 0.955 / 0.739 0.955 / 0.739
Datasets(Gap/Hour/Day) Time Period Nodes
PEMS03_gap&agg 03/2001 - 03/2024 151
PEMS04_gap&agg 06/2002 - 03/2024 822
PEMS05_gap&agg 03/2012 - 03/2024 103
PEMS06_gap&agg 12/2009 - 03/2024 130
PEMS07_gap&agg 06/2002 - 03/2024 3062
PEMS08_gap&agg 03/2001 - 03/2024 212
PEMS10_gap&agg 06/2007 - 03/2024 107
PEMS11_gap&agg 09/2002 - 03/2024 521
PEMS12_gap&agg 01/2002 - 03/2024 1543
tfNSW 01/2013 - 05/2024 27
@@ 5.2. Baselines
In our comparison experiments, we adopted four popular baselines, including MLP, Transformer, and Mamba architectures. Informer (zhou2021informer) introduces an efficient transformer for long sequence time-series forecasting using ProbSparse self-attention and self-attention distilling, enabling encoder-decoder architectures to handle long sequences effectively. MICN (wang2023micn) proposes a multi-scale context network that models both local and global contexts for long-term time series forecasting, capturing patterns across different temporal scales to enhance performance. FEDformer (zhou2022fedformer) introduces a frequency-enhanced decomposed transformer that models time series in both time and frequency domains, improving long-term forecasting by effectively capturing temporal patterns. PatchTST (nietime) applies transformers to time series by treating them as sequences of patches, enabling effective long-term forecasting through self-attention over patch representations to capture temporal dependencies.Autoformer (wu2021autoformer), an earlier SOTA model, leverages a decomposition architecture and auto-correlation mechanism to enhance efficiency and accuracy in long-term time series forecasting, outperforming traditional Transformer models.
iTransformer (liu2024itransformer) is the latest and most effective Transformer-based model, utilizing attention and feed-forward networks on inverted dimensions, embedding time points into variate tokens. DLinear (zeng2023transformers) challenges the effectiveness of Transformer models by proposing a simple one-layer linear model that captures temporal relations in an ordered set of continuous points. It employs positional encoding and uses tokens to embed sub-series, preserving some ordering information in Transformers. Lastly, Mamba (gu2023mamba), a well-known sequential model from last year, uses a bidirectional Mamba block to extract inter-variate correlations and temporal dependencies. Additionally, we have selected five SOTA baselines  (yu2018spatio; guo2019attention; wu2019graph; bai2020adaptive; jiang2023pdformer) from traffic forecasting domain.
Meanwhile, as shown in Table 6, whether using 10% of the data, 50% of the data, or the full dataset, early stopping still occurred within the first or second epoch, and the results remained largely similar. This suggests that the amount of sampled training data does not substantially affect the outcomes, as the task itself is inherently difficult. This further highlights the major challenge introduced in this work: how to effectively leverage the available data to forecast traffic flow far into the future.
@@ 5.3. Implementation Details
We adopted the default settings of the Time-Series-Library (wu2022timesnet) to conduct a comprehensive comparison of baselines. We use the results of five random seeds as the average. We used 96 time steps as input to predict 96, 192, 336 steps. The code was implemented in PyTorch and executed on a V100 GPU with 32GB memory and 384GB RAM, provided by NCI Australia, an NCRIS-enabled capability supported by the Australian Government.
Table 6. Results of different fractions of data in MICN.
Fraction of Data 10% 50% 100%
Metrics MSE MAE MSE MAE MSE MAE
PEMS03_gap 1-year 96 0.514 0.515 0.990 0.736 0.910 0.683
192 0.528 0.513 0.889 0.672 0.814 0.651
336 0.493 0.491 0.808 0.650 0.744 0.612
@@ 5.4. Results of Extremely Long Forecasting with Gaps
We use Mean Squared Error (MSE) and Mean Absolute Error (MAE) metrics to evaluate performance, averaging results across different seeds. It is observed that nearly all results are poor, highlighting the significant challenge posed by domain shifts over time for extremely long forecasting with gaps. These baseline results also indicate that traditional SOTA rankings and methodologies are no longer effective. Notably, MICN, which performs the worst under conventional data and settings, shows the best performance in our setting. This is understandable because MICN’s core technology focuses on exploring the correlation within the data itself. This insight suggests that future efforts to tackle this problem should place greater emphasis on leveraging the intrinsic potential of the data. Additionally, we have compiled the training time for one epoch of all the baselines on the largest dataset, PEMS04_gap, and the smallest dataset, tfNSW, as shown in Figure 6. The reason the PEMS08 dataset yields missing results in Informer is that it is completely unfit for modeling, resulting in extremely poor performance.
Considering that our dataset provides an extensive temporal range for training, we can theoretically extend the input step length considerably. We extended the 96 steps input from Table 3 to a maximum of 1440 steps for testing. As shown in Table 7, the performance improves significantly with the increase in input step length, further demonstrating the substantial potential of our dataset for deep exploration. And the results of datasets PEMS05_gap, PEMS06_gap, PEMS07_gap, PEMS10_gap, PEMS11_gap, PEMS12_gap are shown in Table 4.
@@ 5.5. Results of Hourly and Daily Forecasting
As shown in Table 8, we believe that both the hourly and daily datasets are equally significant. Multi-scale, diverse datasets can provide the community with valuable references. We observe that the performance degrades progressively from the hourly to the daily to the gap datasets. Smaller time scales help reduce complexity and uncertainty, thereby improving the accuracy of prediction. Research has shown that clustering at different scales can improve model performance (wang2024timemixer). Therefore, our aggregated version of the data will contribute new external features to the community.
@@
Here we also provided the results of District 5,6,10,11,12, as shown in Table 9 and Table 4. The reason why some datasets yield missing results is that they are completely unfit for modeling, resulting in extremely poor performance.
Table 9. Comparison in hourly and daily datasets in District 05,06,10,11,12. The symbol - indicates that the result is an outlier.
@@
5.6. Training Efficiency Analysis
We have provided Figure 6 that compares training times per epoch (in seconds) for various models on PEMS04_gap1 and tfNSW_gap1 datasets. DLinear is the fastest $237.9\text{s}$ and $7.4\text{s}$, while AGCRN is the slowest $13{,}369.1\text{s}$ and $658.6\text{s}$. Mamba shows competitive efficiency $396.3\text{s}$ and $13.7\text{s}$, while PatchTST and FEDFormer have long training times. Overall, DLinear is the most efficient, while AGCRN and PDFormer are computationally intensive.
Figure 6. Model training time comparison. The training time for all baselines per epoch is measured in seconds.
6. Conclusion
In this work, we introduced XXLTraffic, the largest and longest‐spanning public traffic dataset to date, covering over 23 years of sensor data from both Los Angeles (PeMS) and New South Wales (TfNSW). By selecting consistently active sensors and providing both hourly and daily aggregations, we deliver a versatile resource that reflects continuous spatial growth and evolving temporal distributions. To support forecasting “beyond test adaptation,” we curated specialized gap‐based subsets (with 1, 1.5 and 2 year intervals) and established benchmark configurations that expose models to real‐world domain shifts, alongside traditional short- and long-term prediction settings. Our dataset release—including raw data, metadata, preprocessing code, and processed aggregates—offers the community a robust platform for developing and evaluating models under a range of practical constraints and forecasting horizons.
While the unprecedented scale and richness of XXLTraffic unlocks new research directions in spatio-temporal modeling and large-model pretraining, its sheer size does impose significant computational demands. Future work should explore efficient data-subsampling strategies, scalable training algorithms, and integration with emerging large-language and foundation models to fully leverage the dataset’s potential without prohibitive resource requirements.
7. Acknowledgments
We would like to acknowledge the support of Cisco’s National Industry Innovation Network (NIIN) Research Chair Program and the ARC Centre of Excellence for Automated Decision-Making and Society (CE200100005). We acknowledge the resources and services from the National Computational Infrastructure (NCI), which is supported by the Australian Government.